OpenStudy (anonymous):

How do I find dy/dx in terms of x and y of the equation x^(5)+y^(3)+2xy^(2)=33?

OpenStudy (anonymous):

jts differentiate it,answer will come itself

OpenStudy (anonymous):

\[5x^4+3y^2y'+2(y^2+2xy')=0\] solve for \(y'\)

OpenStudy (anonymous):

How did you get \[2(y ^{2}+2xy')?\]

OpenStudy (anonymous):

product rule and of course the chain rule i can explain further if you like

OpenStudy (anonymous):

oh no i can't because i made a mistake good catch

OpenStudy (anonymous):

should be \(2(y ^{2}+2xyy')\)

OpenStudy (anonymous):

Okay, that's what I got, but how do you know where to put y'?

OpenStudy (anonymous):

you are thinking of \(y\) as some (unspecified) function of \(x\) i.e. \(y=y(x)\)

OpenStudy (anonymous):

but suppose for a moment that we knew \(y=\sin(x)\) then how would you take the derivative of \[2x\sin^2(x)?\] you need the product rule and the chain rule

OpenStudy (anonymous):

you would get \[2(\sin^2(x)+2x\sin(x)\cos(x))\]

OpenStudy (anonymous):

we don't know what \(y\) is, but if i replace \(\sin(x)\) by \(y\) in the above, i get \[2(y^2+2xyy')\]

OpenStudy (anonymous):

So I solved for y'..does it look right? \[\frac{ (-5x ^{4} + 2y ^{2} ) }{ (3y ^{2} + 4xy) }\]

OpenStudy (anonymous):

i don't know i didn't actually do it

OpenStudy (anonymous):

yeah that looks good